OpenStudy (k4n3k1_k3n):

For all I know, someone will get this in seconds (Unlike me). I haven't been able to easily learn this so don't expect me to catch on immediately please! Which ordered pair is a solution to the inequality? 4x + y > - 6 A. (1, -12) B. (0, -9) C. (-1, -1) D. (-3, 0)

OpenStudy (retireed):

Okay thanks for the heads up, I will try to be gentle.

OpenStudy (retireed):

All you need to do is plug in the values, do the arithmetic and see if the left hand side of the equation is greater than the right hand side. Left side Right side 4x + y > - 6 x y 4 times 1 plus negative 12 equals what? This is the left hand side Is the answer greater than -6 ? Can you answer the TWO questions listed above?

OpenStudy (k4n3k1_k3n):

Okay, so, based on how you laid it out, what I am supposed to do for this situation is plug in the solutions of each answer to test which matches the equation above, so I would say C.

OpenStudy (retireed):

I plugged the (X,Y) of the first answer option into the inequality. 4x + y > - 6 ????? A. (1, -12) 4(1) + (-12) > -6 4 - 12 = ______ > -6 B. (0, -9) C. (-1, -1) D. (-3, 0)

OpenStudy (retireed):

4x + y > - 6 ????? A. (1, -12) 4(1) + (-12) > -6 4 - 12 = ______ > -6 B. (0, -9) C. (-1, -1) 4(-1) + (-1) > -6 -4 - 1 = -5 > -6 YES -5 is greater than -6 D. (-3, 0)

OpenStudy (k4n3k1_k3n):

Okay! Well thank you for the assistance!